There are 22 students on the debate team. The number of boys is two more than three times the number of girls. How many boys are on the team?
Found 2 solutions by ankor@dixie-net.com, greenestamps:
Answer by ankor@dixie-net.com(22740)

b = no. of boys
g = no. of girls
:
Write an equation for each statement
:
There are 22 students on the debate team.
b + g = 22
or
g = 22-b
The number of boys is two more than three times the number of girls.
b = 3g + 2
How many boys are on the team?
Replace g with (22-b) in the above equation
b = 3(22-b) + 2
b = 66 - 3b + 2
b + 3b = 66 + 2
4b = 68
b = 68/4
b = 17 boys
:
You can find the no. of girls and check this for yourself
:



Answer by greenestamps(13200)



If a solution using formal algebra is not required, find the answer using logical reasoning and a bit of mental arithmetic.

(1) Have the 2 "extra" boys step out of the room for a moment.
(2) That leaves 20 team members, with 3 times as many boys as girls.
(3) That means 15 boys and 5 girls.
(4) Bring the 2 extra boys back in the room, making the answer 17 boys and 5 girls.